OpenStudy (selena2345):

To the nearest tenth,what is the distance between (7,4) and (12,5)? A: 4.8 Units B:4.9 Units C:5.0 Units D: 5.1 Units

OpenStudy (accessdenied):

Have you applied the distance formula in this case?

OpenStudy (anonymous):

i think it's B

OpenStudy (accessdenied):

Why do you think it is B?

OpenStudy (anonymous):

bc i jus do dang

OpenStudy (accessdenied):

If you are just guessing, that is not helpful to the asker... As I said, we should apply the distance formula with the two points: \( d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \) And then use a calculator to determine the approximate form with rounding .

OpenStudy (anonymous):

k bye

OpenStudy (selena2345):

can you show me how ?

OpenStudy (selena2345):

bye @_mar5iah_

OpenStudy (anonymous):

Do what @AccessDenied said and plug in your \((x_1,y_1)\wedge(x_2,y_2)\) into that formula.

OpenStudy (accessdenied):

Just substitute our points into the equation: \( (\color{blue}{12}, \color{blue}{5}) \text{ and } (\color{green}7, \color{green}4) \) \( d = \sqrt{(\color{blue}{12} - \color{green}{7})^2 +( \color{blue}5 - \color{green}4)^2} \) And simplify. That makes sense?

OpenStudy (anonymous):



OpenStudy (selena2345):

i got 626 @AccessDenied

OpenStudy (selena2345):

am i right?@AccessDenied

OpenStudy (selena2345):

@AccessDenied

OpenStudy (accessdenied):

That seems much too high. Could you write out the operations you did in a little more detail so we can figure out the issue? :)

OpenStudy (selena2345):

25^2 + 1^2 d=625+1 d=626

OpenStudy (selena2345):

@AccessDenied

OpenStudy (accessdenied):

I think you squared the 5 twice: (12 - 7)^2 = 5^2 = 25

OpenStudy (selena2345):

ok @AccessDenied

OpenStudy (accessdenied):

sqrt(5^2 + 1^2) would look better now, which simplifies down to: sqrt(26)

OpenStudy (selena2345):

what do i do ? Can u help me with that part ? I Am Confused @AccessDenied

OpenStudy (accessdenied):

Which part? The sqrt(26) ?

OpenStudy (selena2345):

what is the answer ?

OpenStudy (accessdenied):

I agreed with the steps up to here: sqrt( (12 - 7)^2 + (5- 4)^2 ) = sqrt( 5^2 + 1^2 ) = sqrt(26) The answer comes from simply using a calculator now. Enter the sqrt(26) and you should get a decimal. Round it to the nearest tenth.

OpenStudy (selena2345):

26 is rounded to 5.0 units Answer Choice : C Am i right ? c:5.0 units @AccessDenied

OpenStudy (accessdenied):

You originally had 5.099019... ? Since 5.09 is closer to 5.1 than 5.0, so we round up.

OpenStudy (selena2345):

am i right @AccessDenied

OpenStudy (accessdenied):

You are rounding: \( 5.0\underline{\color{red}9}9019... \) We use this digit to determine whether to round up or down. If the digit before the tenths place is 5 or higher, we round UP to 5.1. If the digit before the tenths place is less than 5, we round DOWN to 5.0. So the choice of (C) 5.0 is false because we should be rounding up, or (D) 5.1.

OpenStudy (selena2345):

can you help me with another question ? @AccessDenied